### v1

@@ 4 Methodology
Figure 3: Pipeline of our M3PT. In this method, two modal images of the same size go through five steps: embedding, uni-modal exploration, middle fusion, fusion-modal enhancing, and state estimation, to obtain the predicted bounding box. Here, the L transformer encoder blocks from the upstream model are divided into two groups, for the uni-modal exploration and fusion-modal enhancing modeling stages respectively. The subscripts and superscripts of the modules and symbols indicate the modality and the layer number respectively.
@@ 4.1 Overview
In this work, we propose a novel RGB-T prompt tracking method M3PT that leverages a middle fusion tracking framework and multi-form, multi-stage visual prompts. The overall pipeline of our method is shown in Figure 3.
@@
We first developed a multi-modal tracking framework based on middle fusion. The framework’s backbone consists of two parts for different stages. The first-stage backbone is a dual-stream network that extracts two uni-modal features separately. The second-stage backbone is a single-stream network that enhances the fusion-modal features. The middle fusion module is inserted between the two backbone stages.
@@
Furthermore, we embed the pre-trained foundation model into our framework by using four flexible prompt strategies to achieve effective knowledge transfer. These strategies are: (1) The L transformer encoder blocks from the foundation model backbone are divided into two groups, containing N and M encoder blocks respectively, for feature extraction in two stages. (2) The Uni-modal Exploration Prompt Strategy extends the first-group encoder blocks into a parameter-sharing dual-branch structure, and configures the designed UEP in parallel with them layer by layer, to explore the modality-independent information and generate intra-modal and inter-modal prompts, thus better adapting the encoder blocks to uni-modal feature modeling. (3) The Middle Fusion Prompt Strategy inserts the designed MFP after the first-stage backbone, and feeds the fusion-modal features output by MFP as visual prompts to the backbone in the later stage, to adapt it to the fusion-modal modeling. (4) The Fusion-modal Enhancing Prompt Strategy configures a lightweight prompter FEP in parallel with the second-stage backbone layer by layer, to obtain a richer fusion-modal feature representation and enhance the fusion-modal features propagated forward in the backbone. (5) The Modality-aware and Stage-aware Prompt Strategy prepends three learnable prompts to the input of the backbones in two stages, respectively, to guide them to identify the distribution characteristics of the current modality faster. (6) The head network of the foundation model is directly used to estimate the target state on the search features.
4.2 Uni-modal Exploration Prompt Strategy
@@
Figure 4: The pipeline of our Uni-modal Exploration Prompt Strategy and overall architecture of our designed Uni-modal Exploration-assisted Prompter (UEP). The modality-independent information extracted by the UEP is firstly added to the output of the encoder block of the same modality as intra-modal prompts, and the prompted template tokens are injected as inter-modal prompts in an asymmetric cross-layer manner.
Previous prompt tracking methods hardly exploited any modality-independent information, resulting in insufficient mining and utilization of the discriminative patterns within the modality. To address this issue, we propose a Uni-modal Exploration Prompt Strategy and design a lightweight Uni-modal Exploration-assisted Prompter UEP, to prompt the first-stage backbone to explore and model the uni-modal features.
The proposed UEP is shown in Figure 4. The input tokens $E_{X}$ are first reshaped into features and mapped to a low-dimensional feature space. Then, we configure parallel 1$\times$1 convolutional layer and DW convolutional layer to explore modality-independent channel correlation features through different channel combinations, and subsequently use a 1$\times$1 convolutional layer with activation function GELU to achieve local feature extraction. Finally, the features are remapped to the high-dimensional space. In addition, UEP configures residual connection. Since the feature extraction process is all performed in the low-dimensional space, each UEP contains only a small number of learnable parameters. In our case, we set the dimensions of the low-dimensional feature space of the UEP for the thermal infrared and visible modalities to 8 and 16, respectively, to explore the modality-independent feature representation of the two modalities as much as possible.
@@
Furthermore, to utilize the information to enhance intra-modal discriminative features and achieve effective inter-modal interaction, we generate two kinds of visual prompts based on the output of UEP: intra-modal prompts and inter-modal prompts. Specifically, (1) we add the modality-independent features extracted by UEP as intra-modal prompts to the modality-shared features output by Encoder Block, thus providing a richer uni-modal representation for the next layer of backbone, as shown in Figure 4, taking the visible branch as an example, this hint process can be expressed as:
@@
where subscript $V$ denotes visible modality, superscript $n$ denotes layer number, $H$ denotes input tokens of an encoder block. (2) We separate the template tokens from the prompted modal features $E_{V}^{n}]\in R^{(N_{Z}+N_{X})\times D}$, and use them as inter-modal prompts to add element-wise with the template of another modality, as shown in Figure 4. We adopt an asymmetric cross-layer manner to inject the prompts, specifically, we inject the inter-modal prompts from the thermal infrared modality into the input of visible modality branch at the same layer, and inject the intra-modal prompts from visible modality into the input of thermal infrared modality branch at the next layer:
$[E_{V,Z}^{n},E_{V,X}^{n}]$ $=unconcaten(E_{V}^{n})$ (6)
$P_{V,inter}^{n}$ $=E_{V,Z}^{n}$ (7)
$H_{T,Z}^{n+1}$ $=E_{T,Z}^{n}+P_{V,inter}^{n}$ (8)
$[E_{V,Z}^{n},E_{V,X}^{n}]$ $=unconcaten(E_{V}^{n})$ (9)
$P_{V,inter}^{n}$ $=E_{V,Z}^{n}$ (10)
$H_{T,Z}^{n+1}$ $=E_{T,Z}^{n}+P_{V,inter}^{n}$ (11)
where $unconcaten()$ refers unconcatenation. The reason for using template features as inter-modal prompts is that the templates contain less background noise and more clues about the target appearance. Therefore, by using inter-modal prompts, we provide the upstream backbone with more modality-complementary target appearance clues, which enables the feature extraction of upstream backbone to focus more on the potential regions of the target.
@@ 4.3 Middle Fusion Prompt Strategy
To achieve adaptive fusion of the uni-modal features extracted by the first-stage backbone and provide reliable task prompts for the second-stage backbone, we propose a Middle Fusion Prompt Strategy that utilizes our designed middle fusion prompter MFP to perform adaptive selection of discriminative features and complementary fusion, and views the fusion-modal features as natural visual prompts for the next stage.
MFP is shown in Figure 5. The input feature tokens of the two modalities $E_{V}$ and $E_{T}$ are first reshaped into features FV and FT, and then mapped to low-dimensional feature space by two weight-unshared 1×1 convolutional layers, respectively. Here, we set the dimensions of the low-dimensional feature space of the MFP to 16. Then, inspired by [42], we separate the modality-shared and modality-independent information extracted at each element:
@@
where $F_{V,L}$ and $F_{T,L}$ represent the visible light and infrared features mapped to the low-dimensional space, respectively, $F_{S}^{{}^{\prime}}$ represents the modality-shared features, F’A represents the overall features, $F_{V}^{{}^{\prime}}$ and $F_{T}^{{}^{\prime}}$ represent the visible light and infrared modality-independent features, respectively, $\times$, $+$, and -- represent element-wise multiplication, addition, and subtraction, respectively, and sigmoid operation is used to enhance the local modality-independent features. After that, we concatenate these features in the following order and generate adaptive discriminative filtering weights based on them:
@@
It is noteworthy that we add the shared features $F^{\prime}_{S}$ on top of the weighted fusion, considering that the useful information that may exist at the low-weight positions in the discriminative fusion process will be suppressed, and these low-weight information mostly belong to the modality-shared information. Previous weighted fusion methods have ignored it. Therefore, we ensure the integrity of the effective information by adding $F^{\prime}_{S}$. Finally, the fused features are remapped to the original dimension and reshaped into tokens.
We propose a Middle Fusion Prompt Strategy that inserts MFP after the first-stage backbone, which contains N Encoder Blocks. Unlike [16] and [17], which only perform image-level weighted fusion of the two modal features, we perform adaptive and discriminative selection and complementary fusion of the two modal features that have been fully explored and modeled, and we retain as much useful information as possible. Therefore, the obtained fusion-modal features can serve as reliable fusion-modal prompts injected into the second-stage upstream backbone, making it better adapt to the fusion-modal modeling task.
To the best of our knowledge, this is the first time that the middle fusion framework is introduced to the RGB-T tracking task. Compared with the tracking framework based on image-level fusion, our framework can more effectively explore the independent information of the two modalities; compared with the tracking framework based on feature-level fusion and decision-level fusion, our method can effectively compress the redundant modeling computation and reduce the model parameters. While keeping the total amount of backbone parameters from the pre-trained model unchanged, by flexibly adjusting the proportion allocation of the two-stage backbone parameters, our method can achieve the optimal balance between performance and efficiency.
@@ 4.5 Modality-aware And Stage-aware Prompt Strategy
We intuitively assume that each modality contains some fixed patterns such as low-level distribution characteristics, which can be learned and stored to guide the backbone to quickly and familiarize themselves with the current modality and modeling task. Therefore, we propose a Modality-aware and Stage-aware Prompt Strategy that uses three learnable prompts to learn and store the fixed patterns of the visible light, infrared, and fusion modalities. Inspired by [36], we prepend these prompts to the input of corresponding encoder blocks, to provide clear modality and stage identification for the upstream backbone, as shown in Figure 3.
These prompts only participate in the feature extraction of upstream backbone, and do not participate in the generation and other prompting processes. In other words, these prompts are separated from the total tokens at the input of all prompter branches, and are concatenated with the prompted feature tokens after completing the other prompting processes. Specifically, after MFP completes the fusion of the two modal features, the two separated uni-modal prompts are added to the fusion-modal prompts, and prepended to the input of the second-stage backbone. Since they do not contain positional encoding, these prompts can be inserted into any position of the feature tokens. In our case, we set the token number of the learnable prompts to 2.
@@ 5 Experiment
Figure 6: The evaluation curves of the three metrics PR, SR, and NPR on LasHer.
Figure 7: The evaluation curves of the MPR metric for the 19 annotated challenge attributes on LasHer.
Figure 8: The evaluation curves of the two metrics MPR and MSR on RGBT234.
Figure 9: The evaluation curves of the MPR metric for the 12 annotated challenge attributes on RGBT234.
Figure 10: The evaluation curves of the two metrics PR and SR on RGBT210.
Figure 11: The evaluation curves of the two metrics MSR and MPR on the long-term test subset of VTUAV.
Figure 12: The evaluation curves of the two metrics MSR and MPR on the short-term test subset of VTUAV.
@@
5.1 Experimental Settings
5.1.1 Implementation Details
We use the pre-trained OSTrack-256 model as the foundation of our method. We fine-tune the prompt-related parameters that we designed on the LasHer’s train split and freeze all the other parameters of the foundation model. We set the batch size to 32 and the number of epochs to 60. We use the AdamW optimizer with a weight decay rate of 0.0001 and an initial learning rate of 0.0004, which decays to 0.1 times the original value after 48 epochs. The training process is end-to-end. After the offline training, we do not perform any online parameter update operations which may reduce our model’s inference speed. The template image size is set to 128×128 and the search image size is set to 256×256. The block number of the first-stage backbone is set to 2. We conduct all experiments on four Nvidia RTX 3090 GPUs, using Python and PyTorch.
5.1.2 Experimental Benchmarks
We evaluate the performance of our method on four public and challenging RGBT tracking benchmarks: LasHer [35], RGBT234 [43], RGBT210 [44], and VTUAV [15]. We will briefly describe them in the following paragraphs.
LasHer: LasHer is the largest-scale RGB-T tracking benchmark, containing 1224 highly aligned video sequence pairs and 734.8K image pairs. The benchmark covers 32 common object categories and 19 typical challenge attributes, posing great challenges for current RGB-T tracking algorithms.
RGBT210: RGBT210 consists of 210 video sequence pairs and 210K image pairs, with the longest video sequence having 8K frames.
RGBT234: RGBT234 is the extension of RGBT210, containing 234 RGBT video sequence pairs and 234K image pairs. The benchmark includes 12 common challenge attributes.
VTUAV: VTUAV is a large-scale benchmark, comprising 500 video sequence pairs and 1700K image pairs, with high resolution. It is divided into long-term subset and short-term subset, each containing 250 video sequences. VTUAV is the first RGB-T tracking benchmark that formally proposes the long-term tracking task.

### v2

@@ 4 Methodology
Figure 3: Pipeline of our M3PT. In this method, two modal images of the same size go through five steps: embedding, uni-modal and inter-modal exploration, middle fusion, fusion-modal enhancing, and state estimation, to obtain the predicted bounding box. Here, the L transformer encoder blocks from the upstream model are divided into two groups, for the uni-modal and inter-modal exploration and fusion-modal enhancing modeling stages respectively. The subscripts and superscripts of the modules and symbols indicate the modality and the layer number respectively.
@@ 4.1 Overview
In this work, we propose a novel RGB-T prompt tracking method M3PT that leverages our designed middle fusion meta-framework and multi-form, multi-stage visual prompts. The overall pipeline of our method is shown in Figure 3.
@@
We first develop a universal middle fusion meta-framework for multi-modal tracking task. In the meta-framework, the tracker’s backbone consists of two parts for different stages: The first-stage backbone is a dual-stream network that extracts two uni-modal features separately; The second-stage backbone is a single-stream network that enhances the fusion-modal features. The middle fusion module is inserted between the two backbone stages.
Furthermore, based on the meta-framework, we embed the pre-trained foundation model into our tracking pipeline by using four flexible prompt strategies to achieve effective knowledge transfer. These strategies include Uni-modal and Inter-modal Exploration Prompt Strategy, Middle Fusion Prompt Strategy, Fusion-modal Enhancing Prompt Strategy and Modality-aware and Stage-aware Prompt Strategy.
@@
Therefore, the comprehensive pipeline of our prompt tracking approach is delineated below: (1) Following the meta-framework, the L transformer encoder blocks from the foundation model backbone are divided into two groups, containing N and M encoder blocks respectively, for feature extraction in two stages. (2) The Uni-modal and Inter-modal Exploration Prompt Strategy extends the first-group encoder blocks into a parameter-sharing dual-branch structure, and configures our designed UEP and IP in parallel with them layer by layer, to explore the modality-independent information and operate scenario-adaptive cross-modal communication, thus better adapting the encoder blocks to modeling uni-modal clues sufficiently in various intricate scenarios. (3) The Middle Fusion Prompt Strategy inserts the designed MFP after the first-stage backbone, and feeds the fusion-modal features output by MFP as visual prompts to the backbone in the later stage, to adapt it to the fusion-modal modeling. (4) The Fusion-modal Enhancing Prompt Strategy configures a lightweight prompter FEP in parallel with the second-stage backbone layer by layer, to obtain a richer fusion-modal feature representation and enhance the fusion-modal features propagated forward in the backbone. (5) The Modality-aware and Stage-aware Prompt Strategy prepends three kinds of learnable modality and stage prompts to the input of the backbones in two stages, respectively, to guide foundation model to identify the distribution characteristics of the current modality faster. (6) The head of foundation model is frozen and directly utilized to estimate the target state on the search features.
4.2 Uni-modal and Inter-modal Exploration Prompt Strategy
Previous prompt tracking methods, like ProTrack [27], ViPT [28] and OneTracker [30], hardly exploit enough modality-independent information and also fail to facilitate priority-aware inter-modal interaction across diverse uncontrolled environments. To tackle these problems, we propose the Uni-modal and Inter-modal Exploration Prompt Strategy, formulating two lightweight components: Uni-modal Exploration-assisted Prompter (UEP) and Inter-modal Self-adaptive Prompter (IP). Employing our prompt strategy and prompters, the first-stage backbone is directed to explore the uni-modal features while achieving scenario-adaptive inter-modal communication. The pipeline of our strategy and overall architechure of UEP and IP are shown in Figure 4.
@@ 4.2 Uni-modal and Inter-modal Exploration Prompt Strategy
Figure 4: The pipeline of our Uni-modal and Inter-modal Exploration Prompt Strategy and overall architecture of two lightweight prompters which include Uni-modal Exploration-assisted Prompter (UEP) and Inter-modal Self-adaptive Prompter (IP). In the prompt strategy, the modality-independent information extracted by our designed UEP is firstly added to the output of the encoder block of the same modality as intra-modal prompts, and the prompted template tokens are further utilized by our designed IP to generate effective inter-modal scenary prompts.
Uni-modal Exploration-assisted Prompter (UEP). We design UEP for sufficient mining the discriminative patterns within each modality. As shown in Figure 4, the input tokens $E_{X}$ ($X\in\left\{V,T\right\}$) are firstly reshaped into features $F_{X}$ and mapped to a low-dimension feature space, whose purpose is reducing redundant features and trainable parameters. In the low-dimension space, we configure parallel standard 1$\times$1 convolutional layer, depth-wise convolutional layer and dilated convolutional layer to capture standard features, channel-private features, and larger receptive field local features, respectively. Through this way, modality-independent information can be thoroughly explored from various perspectives. Subsequently, an additional standard 1$\times$1 convolutional layer with activation function GELU is employed to integrate above features. Assuming extract modality-independent features of visible modality, the entire process conducted in the low-dimensional space can be succinctly expressed as follows:
$F_{sum}$ $=Cv_{1}(F_{V,L})+Cv_{DL}(F_{V,L})+Cv_{DW}(F_{V,L})$ (4)
$F^{\prime}_{V,L}=GELU(Cv_{2}(F_{sum}))$ (5)
Where subscripts $L$ and $V$ respectively denote low dimension and visible modality, $Cv_{1}()$ and $Cv_{2}()$ all denote standard 1$\times$1 convolutional layer, $Cv_{DW}()$ and $Cv_{DL}()$ respectively denote depth-wise convolutional layer and dilated convolutional layer, and $F^{\prime}$ denotes integrated features. Finally, the combined features are remapped to original high-dimension space. In addition, UEP configures residual connection. Since the feature extraction process is all performed in the low-dimension space, each UEP contains only a small number of learnable parameters. In our case, the dimensions of the low-dimension feature space of UEP is set to 8.
Inter-modal Self-adaptive Prompter (IP). Our Inter-modal Self-adaptive Prompter is designed to adapt the foundation model to perform robustly in priority-agnostic intricate environment. As shown in Figure 4, the template tokens of prompting and prompted modality are firstly reshaped into features $F_{P,Z},F_{X,Z}$ (Here $P$ and $X$ respectively denote prompting and prompted modality, and $Z$ denotes template.) and mapped to a low-dimension feature space. The reason for using template features as inter-modal prompts is that the templates contain less background noise and more clues about the target appearance. Then, we employ average channel-pooling to capture cross-channel clues of the prompting modality features $F_{P,Z,L}$ ($L$ denotes low dimension.) and take the one-channel feature as a channel-level inter-modal prompt $P_{P,inter}$. To elaborate, we append the prompt $P_{P,inter}$ as an as an auxiliary channel to the prompted features $F_{X,Z,L}$ and amalgamate them via a standard 1$\times$1 convolution layer. Assuming we designate visible modality as the prompting modality and thermal infrared modality as the prompted modality, the entire process conducted in the low-dimensional space can be succinctly expressed as follows:
$P_{V,inter}=ACP(F_{V,Z,L})$ (6)
$F^{\prime}_{T,Z,L}$ $=Cv(concaten([P_{V,inter},F_{T,Z,L}]))$ (7)
where $ACP()$ represents average channel pooling, $Cv()$ denotes a standard 1$\times$1 convolution layer and $F^{\prime}$ denotes combined prompted features. Finally, a high-dimension projection and a residual connection are configured, which is similar to UEP. In our case, the dimensions of the low-dimension feature space of IP is set to 8. Why we choose this channel-level prompt approach? It is noteworthy that existing prompt methods primarily sum the two modalities with employ constant global weights or element-wise weights (acquired through learning), to achieve inter-modal communication. While the latter methods generally surpass the former in effectiveness, static weights consistently fall short in adapting the model proficiently to the dynamic modality dominance in complex, open-world scenarios. Different with them, our prompter further learns how to infuse the critical information of the prompting modality, captured via channel pooling, into the prompted modality, rather than indiscriminately applying fixed weights to both modalities. Consequently, our strategy demonstrates enhanced adaptability when confronted with complex scenarios characterized by variable modality priority.
@@
The proposed Uni-modal and Inter-modal Exploration Prompt Strategy utilize modality-private UEP and modality-sharing IP to adapt first stage backbone to thoroughly explore the coupled feature space of the two modalities as illustrated in Figure 4. Specifically, the prompt pipeline could be described as follows: (1) For uni-modal prompt process, the modality-independent features extracted by UEP is added as intra-modal prompts to the modality-shared features output by Encoder Block, thus providing a richer uni-modal representation for the next backbone layer. Taking the visible branch as an example, this prompt process can be expressed as:
@@
where superscript $n$ denotes layer number, $H$ denotes input tokens of an encoder block, and $P$ denotes the prompts generated by designed prompter. (2) For inter-modal prompt process, current and another modality branches are respectively taken as prompting and prompted features, the template tokens of them are input to IP, realizing inter-modal prompting. Notably, here we adopt an asymmetric cross-layer manner for inter-modal prompting, which could formulated as follows:
$H_{T,Z}^{n}$ $=IP^{n}(E_{T,Z}^{n-1},E_{V,Z}^{n})$ (10)
$H_{V,Z}^{n+1}$ $=IP^{n}(E_{T,Z}^{n},E_{V,Z}^{n})$ (11)
@@ 4.3 Middle Fusion Prompt Strategy
To achieve full integration of the uni-modal features extracted by the first-stage backbone and provide reliable task prompts for the subsequent backbone, we propose the Middle Fusion Prompt Strategy that utilizes our designed Middle Fusion Prompter (MFP) to perform active selection of discriminative features and self-adaptive fusion, and take the fusion-modal features as natural visual prompts for the next stage.
Middle Fusion Prompter (MFP). MFP is shown in Figure 5. The input feature tokens of the two modalities $E_{V}$ and $E_{T}$ are first reshaped into features $F_{V}$ and $F_{T}$, and then mapped to low-dimension feature space by two modality-private standard 1$\times$1 convolutional layers, respectively. Here, we set the dimensions of the low-dimension feature space of the MFP to 16. Then, inspired by [23], we separate the modality-shared and modality-independent information extracted at each element:
@@
where $F_{V,L}$ and $F_{T,L}$ represent the visible light and infrared features mapped to the low-dimension space, respectively, $F^{\prime}_{S}$ represents the modality-sharing features, $F^{\prime}_{A}$ represents the overall features, $F^{\prime}_{V}$ and $F^{\prime}_{T}$ represent the visible light and infrared modality-independent features, respectively, $\times$, $+$, and -- represent element-wise multiplication, addition, and subtraction, respectively, and sigmoid operation is used to enhance the local modality-independent features. After that, we concatenate these features in the following order and generate adaptive discriminative filtering weights based on them:
@@
It is noteworthy that we add $F^{\prime}_{S}$ to the weighted fusion features, considering that certain useful information existing at the low-weight positions in the discriminative fusion process will be suppressed and these low-weight information mostly belong to the modality-sharing information. Previous weighted fusion methods have ignored it. Therefore, we ensure the integrity of the effective information by adding $F^{\prime}_{S}$. Finally, the fused features are remapped to the original dimension and reshaped into tokens.
We propose the Middle Fusion Prompt Strategy that inserts MFP after the first-stage backbone, which contains N Encoder Blocks. Unlike [27] and [28], which only perform image-level weighted fusion of the two modal features, we perform adaptive and discriminative selection and complementary fusion of the two modal features that have been fully explored and modeled, and we retain as much useful information as possible. Therefore, the obtained fusion-modal features can serve as reliable fusion-modal prompts injected into the second-stage upstream backbone, making it better adapt to the fusion-modal modeling task.
Furthermore, to the best of our knowledge, this is the first time that the middle fusion meta-framework is introduced to the RGB-T tracking task. Compared to the image-level fusion framework, our meta-framework can help the tracker more effectively explore the independent information of the two modalities; compared to the feature-level fusion and decision-level fusion framework, our meta-framework can help the tracker effectively compress the redundant modeling computation and reduce the model parameters. Our prompt method provide a typical application instance. While keeping the total amount of backbone parameters from the pre-trained model unchanged, by flexibly adjusting the fusion location, the proportion allocation of the two-stage backbone parameters could be freely changed, balancing the performance and efficiency for diverse application requirements.
@@ 4.4 Fusion-modal Enhancing Prompt Strategy
where superscript denotes layer number, $H$ denotes input tokens of an encoder block, and $P$ denotes the prompts generated by designed prompter.
@@ 4.5 Modality-aware And Stage-aware Prompt Strategy
We intuitively assume that each modality contains some fixed patterns such as low-level distribution characteristics, which can be learned and stored to guide the backbone to quickly and familiarize themselves with the current modality and modeling task. Therefore, we propose Modality-aware and Stage-aware Prompt Strategy that stores the fixed patterns of the visible, thermal, and fusion modalities by three kinds of learnable prompts. Inspired by [49], we prepend these prompts to the input of corresponding encoder blocks, to provide upstream backbone with guidance of current modality and stage, as shown in Figure 3.
These prompts only participate in the feature extraction of upstream backbone. Namely, the prompts are separated from total tokens at the input of all prompter branches, and are concatenated with the prompted feature tokens after completing other prompting processes. For an instance, after MFP completes the modality feature fusion, the two uni-modal prompts are added to the fusion-modal prompts which then are prepended to the input of second-stage backbone. Since they do not contain positional encoding, these prompts can be inserted into any position of the feature tokens. In our case, we set the token number of the learnable prompts to 2.
@@ 5 Experiment
Figure 6: The evaluation curves of metrics PR, SR and NPR on LasHer.
Figure 7: The evaluation curves of MPR for 19 challenge attributes on LasHer.
Figure 8: The evaluation curves of metrics MPR and MSR on RGBT234.
Figure 9: The evaluation curves of MPR for 12 challenge attributes on RGBT234.
Figure 10: The evaluation curves of metrics PR and SR on RGBT210.
Figure 11: The evaluation curves of metrics PR and SR on GTOT.
Figure 12: The evaluation curves of metrics MPR and MSR on long-term test subset of VTUAV.
Figure 13: The evaluation curves of metrics MPR and MSR on short-term test subset of VTUAV.
@@
5.1 Implementation Details
Foundation Model. We choose the pre-trained OSTrack-256 model as the foundation model and fine-tune our M3PT in RGB-T tracking task.
Training. For fair evaluation and comparison, we train our model on LasHer training set and whole LasHer, respectively. During training, we freeze the foundation model and only train prompt parameters. The batch size, number of epochs and sample pairs of each epoch are set to 32, 60 and 60000, respectively. We use AdamW optimizer with a weight decay rate of 0.0001 and an initial learning rate of 0.0004, which decays to 0.1 times the original value after 48 epochs. The training process is end-to-end. The template and search image sizes are set to 128×128 and 256×256, respectively. The block number of the first-stage backbone is set to 10. To regulate the number of fine-tuning parameters during the fine-tuning process, we selectively employ UEP only for the 2nd, 5th and 8th Encoders. We conduct our training on four Nvidia RTX 3090 GPUs, using Python and PyTorch.
Inference and Evaluation. The model trained on Lasher training set and whole LasHer are evaluated on LasHer testing set and other benchmarks, respectively. There are not online parameter or template update operated during the inference.
5.2 Experimental Benchmarks
To sufficiently verify the effectiveness of our method, we evaluate its performance on six challenging and public RGB-T tracking benchmarks: LasHer [48], RGBT234 [24], RGBT210 [55], VTUAV [17], GTOT [56] and VOT2019RGB-TIR [57], which encompass a diverse range of tracking scenarios, challenging attributes, and object categories, spanning rich long and short video sequences. We will briefly describe them in the following paragraphs.
LasHer: LasHer is the largest-scale RGB-T tracking benchmark up to now, containing 1224 highly aligned video sequence pairs and 734.8K image pairs. The benchmark covers 32 common object categories and 19 typical challenge attributes, posing great challenges for current RGB-T tracking algorithms. The official evaluation metrics of LasHeR are Precision Rate (PR), Normalized Precision Rate (NPR), and Success Rate (SR).
RGBT210: RGBT210 consists of 210 video sequence pairs and 210K image pairs, with the longest video sequence having 8K frames. The official evaluation metrics of RGBT210 are PR and SR.
RGBT234: RGBT234 is the extension version of RGBT210, containing 234 RGB-T video sequence pairs and 234K image pairs. The benchmark includes 12 common challenge attributes. The official evaluation metrics of RGBT234 are Maximum Precision Rate (MPR) and Maximum Success Rate (MSR).
GTOT: GTOT is the first large-scale benchmark in RGB-T tracking field. 50 pairs of annotated RGB-T sequences are contained in its dataset. PR and SR are employed as the official evaluation metrics of GTOT.
VTUAV: VTUAV is a large-scale benchmark, comprising 500 video sequence pairs and 1700K image pairs, with high resolution. It is divided into long-term subset and short-term subset, each containing 250 RGB-T video sequence pairs. VTUAV is the first RGB-T tracking benchmark that formally proposes the long-term tracking task. The official evaluation metrics employed on VTUAV are MPR and MSR.
VOT2019RGB-TIR: VOT2019RGB-TIR benchmark is introduced in the RGB-T tracking channel of VOT2019 [57] tacking challenge and contains 60 pairs of RGB-T sequences. The evaluation metrics employed on the benchmark are Expected Average Overlap (EAO), Accuracy (AUC), and Robustness (R).